### v1

@@ 2.1 Assumptions for identification
To ensure the identification of the treatment effect, we introduce a series of assumptions about the DGPs, in addition to the assumption regarding the functional form. These assumptions are crucial for achieving consistent estimation of the treatment effect, yet they are easy to satisfy in practice.
@@
    Covariate orthogonality.
@@
    Moment condition of factors and covariates.
@@
    Compact parameter space for $\Gamma$.
The assumptions outlined above serve as regularity conditions necessary for ensuring the consistency of the estimator. The first condition reflects the exclusion restriction commonly associated with instrumental variable regression, indicating that the instruments (in this case, covariates) are orthogonal to the error terms. This orthogonality ensures that the covariates do not correlate with any unobserved error terms influencing the outcome. Similarly, the functional form described in Equation 3 aligns with the framework utilized in two-stage instrumental regression. The second condition imposes a limitation on the second moment of a series of random variables, ensuring that their variances remain finite and do not escalate indefinitely. The third condition mandates that the parameter space for the mapping matrix $\Gamma$ be compact, thereby circumventing issues related to rank deficiency. This compactness is vital for the estimability of $\Gamma$ and the existence of its inverse, mirroring standard factor analysis assumptions. We provide detailed explanations of each assumption’s conditions in Appendix A.2.
@@
Assumptions for asymptotic normality.
@@
    Panel-wise and cross-sectional central limit theorem.
@@
    Cross-sectional homoskedasticity of covariates.
To derive the asymptotic properties of the CSC-IPCA estimator, we introduce additional assumptions. Assumption 3 encompasses panel-wise and cross-sectional central limit theorems for various variables, which are fulfilled by diverse mixing processes. These conditions are pivotal for determining the asymptotic distribution of factors and mapping matrix estimations. For an in-depth discussion, we refer to Kelly et al. (2020). The requirement of bounded dependence stipulates that both time series and cross-sectional dependencies of $X_{it}\epsilon_{it}$ are bounded, a crucial step for establishing asymptotic normality. Meanwhile, the assumption of cross-sectional homoskedasticity simplifies the expressions for the asymptotic variances of the estimated mapping matrix $\Gamma$. It should be noted that relaxing this assumption would not alter the convergence rate. Detailed explanations of each assumption’s conditions are provided in Appendix Assumption A.3.
@@ 3 Estimation
The CSC-IPCA estimator of the treatment effect for a treated unit $i$ at time $t$ is defined as the difference between the observed outcome and its estimated counterfactual: $\hat{\delta}_{it}=Y_{it}^{1}-\hat{Y}_{it}^{0}$. To combine the functional form in Equation 3, we get the structural component of the potential outcome $Y_{it}=(X_{it}\Gamma)F^{\prime}_{t}$. The CSC-IPCA method is estimated by minimizing the sum of squared residuals of the following objective function:
@@
Unlike the IFE method (Bai (2009); Xu (2017)), our approach requires estimating only two parameters, $\Gamma$ and $F_{t}$, simplifying the process. Different from principal component analysis (Jolliffe (2002); Stock and Watson (2002)), our method involves using covariates to instrument the factor loadings. This necessitates the estimation of $\Gamma$ rather than $\Lambda_{i}$, so we can not directly use eigenvalue decomposition. While the objective function in Equation 6 formulates the problem as minimizing a quadratic function with a single unknown variable (e.g., $\Gamma$) while holding the other variable (e.g., $F_{t}$) constant. This structure enables the application of the alternating least squares (ALS) method for a numerical solution. Generally, the imputation for the missing counterfactual $Y_{it}(0)$ is executed in four steps:
@@
Step 2: The subsequent step involves estimating the mapping matrix $\hat{\Gamma}_{treat}$ for treated unit $i$ at time $t$, employing the previously estimated time-varying factors $\hat{F}_{t}$ and the observed covariates $X_{it}$, using treated group data exclusively for the pre-treatment period.
@@
$F_{norm}$ $=R^{-1}\hat{F}_{t},$
$s.t.\Gamma_{norm}^{\prime}\Gamma_{norm}$ $=\mathcal{I}_{K},\quad F_{norm}F_{norm}^{\prime}/T=\text{Diagonal}.$
Similar to most factor analysis methods, the estimated $\Gamma$ and $F_{t}$ are not deterministic. There exist an infinite number of “rotated” parameters $\Gamma R$ and $R^{-1}F_{t}$ that yield the same objective function value (i.e., $X_{it}\Gamma F_{t}=X_{it}\Gamma RR^{-1}F_{t}$). To make the estimation identifiable, it is necessary to impose some constraints on the estimated $\Gamma$ and $F_{t}$. Following Connor and Korajczyk (1993); Stock and Watson (2002); Bai and Ng (2002), the aforementioned restrictions reduce the model’s complexity and make it easier to understand and interpret the relationships between the factors $F_{t}$ and factor loadings $\Lambda_{it}$. Unlike the estimation methods in Bai (2009); Xu (2017), where normalization constraints are set before the estimation, the structural component $X_{it}\Gamma F_{t}$ allows us to normalize the estimated $\Gamma$ and $F_{t}$ after the estimation. This enables us to easily find the rotation matrix $R$ that satisfies the above constraints.
Step 4: The final step involves imputing the counterfactual outcome $\hat{Y}_{it}^{0}$ for treated unit $i$ at time $t$ by substituting the estimated mapping matrix $\hat{\Gamma}_{norm}$ and the time-varying factors $\hat{F}_{norm}$ into the following equation:
$\hat{Y}_{it}(0)=(X_{it}\hat{\Gamma}_{norm})\hat{F}^{\prime}_{norm},\quad\forall i\in\mathcal{T}\ \&\ T_{pre}<t\leq T.$ (10)
The main difference between CSC-IPCA and the instrumented principal component analysis (IPCA) as proposed by Kelly et al. (2020) lies in the purpose of prediction. In the IPCA method, the authors predict the next period stock returns using all covariates from the preceding period, under the assumption that the mapping matrix $\Gamma$ remains constant across all observations. In contrast, CSC-IPCA introduces a pivotal distinction: it operates under the assumption that treated and control groups are characterized by unique mapping matrices, $\Gamma_{\text{treat}}$ and $\Gamma_{\text{ctrl}}$. This assumption is vital for the unbiased estimation of the ATT, setting the CSC-IPCA method apart by directly addressing heterogeneity in the treatment effect through the specification of group-specific mapping matrices. The detailed estimation procedures are presented in the Appendix A.3.
@@ 3.2 Inference
In the context of causal inference, the pioneering application of factor models is attributed to Hsiao et al. (2012). Nonetheless, a formal framework for inference using this method was not established until the works of Chan et al. (2016) and Li (2018). These methods of inference rely on a large number of control units and pre-treatment periods to develop asymptotic properties. In recent years, conformal inference (Chernozhukov et al. (2021)) has gained popularity in the causal inference literature, as seen in Ben-Michael et al. (2021), Roth et al. (2023), and Imbens (2024). Conformal inference is a nonparametric method that provides exact and robust inference without requiring the specification of the model. Our causal inference framework, designed for predicting missing counterfactuals, allows us to construct inference procedures based on conformal prediction, as introduced by Shafer and Vovk (2008), to ensure robustness against misspecification. The causal effect is identified as the difference between the observed outcomes and these estimated counterfactuals, expressed mathematically as:
$\theta_{it}=Y^{1}_{it}-\hat{Y}^{0}_{it},\quad\forall i\in N_{treat}\ \&\ T_{pre}+1\leq t\leq T,$
where $\theta_{it}$ denotes the treatment effect for unit $i$ at time $t$, and $\hat{Y}_{it}$ represents the imputed counterfactual outcome.
@@ 4 Monte Carlo Simulation
In this section, we employ Monte Carlo simulations to assess the performance of the CSC-IPCA estimator in finite sample settings. We compare the CSC-IPCA estimator against the CSC-IFE and the original SCM estimators. Our comparative analysis focuses on key metrics, including bias, mean squared errors, and convergence properties.
We initiate our analysis with a data generating process that incorporates $L=10$ covariates and $K=3$ time-varying common factors, along with unit and time-fixed effects:
@@
where $X_{it}=[x_{it}^{1},\ldots,x_{it}^{L}]$ denotes a vector of $L\times 1$ covariates, which follows a VAR(1) process. $X_{it}=\mu_{i}+A_{i}X_{i,t-1}+\nu_{it}$, where $A_{i}$ is a $L\times L$ variance-covariance matrix [^6] In our methodology, the variance-covariance matrix is not constrained to be diagonal, thus allowing covariates within each unit to be correlated, reflecting the typical scenario in most economic time series data. To emphasize the independence among different units, we generate $N$ unique variance-covariance matrices, each corresponding to a unit, ensuring cross-sectional independence and preserving time-series correlation. Moreover, we impose a condition on these matrices by requiring the eigenvalues of $A_{i}$ to have characteristic roots that reside inside the unit circle, thereby assuring the stationarity of the VAR(1) process., The drift term $\mu_{i}$ equals 0 for control units and 2 for treated units [^7] This configuration underscores that the treatment assignment is not random; rather, it depends on the covariates $X_{it}$., and $\nu_{it}$ is a $L\times 1$ vector of i.i.d. standard normal errors. While $F_{t}=[f_{t}^{1},\ldots,f_{t}^{K}]$ denotes the vector of common factors, adhering to a similar VAR(1) process, the variable $\epsilon_{it}$ represents the idiosyncratic error term. Unit and time fixed effects, $\alpha_{i}$ and $\xi_{t}$ respectively, are uniformly drawn from the interval $(0,1)$. The coefficient vector $\beta=[\beta^{1},\ldots,\beta^{L}]$ associated with the covariates is drawn uniformly from $(0,1)$, and $\Gamma$, the $L\times K$ mapping matrix for the factor loadings, is drawn uniformly from $(-0.1,0.1)$. The treatment indicator $D_{it}$ is binary, defined as $D_{it}=1$ for treated units during post-treatment periods, and $D_{it}=0$ otherwise. The heterogeneous treatment effect is modeled as $\delta_{it}=\bar{\delta}_{it}+e_{it}$, where $e_{it}$ is i.i.d as standard normal, and $\bar{\delta_{t}}=[0,\cdots,0,1,2,\ldots,T_{post}]$ represents a time-varying treatment effect [^8] Here we simplify the treatment effect to be constant across units, however the heterogeneous treatment effect across units can also be easily employed.. Only the outcome $Y_{it}$, the covariates $X_{it}$, and the treatment indicator $D_{it}$ are observed, while all other variables remain unobserved.
@@
In this graphic, the upper panel plots simulated data following the above data generating process. The light blue lines represent treated units and the light gray lines represent controls. Key parameters are $N_{treat}=5,N_{ctrl}=45,T_{0}=20,T_{1}=10,L=10$. The lower panel plots a simple event study.
Figure 1 represents the simulated data following our data generating process. Observations from the upper panel indicate that the parallel trend assumption is not met. To verify this, we plot a simple event study, clearly revealing a failure in the parallel trend assumption. Furthermore, outcomes for treated units are marginally higher than for control units. In such cases, the synthetic control method will be biased, as it avoids extrapolation and typically fits poorly for treated units.
@@ 4.1 A simulated example
Following this data generating process, Figure 2 illustrates both the raw data and the imputed counterfactual outcomes as estimated by the CSC-IPCA method. In the upper panel, control units are represented in gray and treated units in light blue, with the average outcome for treated units highlighted in orange. The imputed synthetic average for treated outcomes is also shown, delineated by an orange dashed line. The CSC-IPCA method is capable of capturing the trajectory of the average outcome for treated units before treatment, and we observe the divergence after the treatment. The lower panel of Figure 2 shows the estimated ATT (dashed line) alongside the true ATT (solid line) and the 95% confidence interval based on conformal inference. The CSC-IPCA method is able to capture the true ATT, as evidenced by the close alignment between the dashed and solid lines. The confidence interval constructed through conformal inference is also accurate, as it encompasses the true ATT.
@@ 4.3 Finite sample properties
We present the Monte Carlo simulation results in Table 1 to investigate the finite sample properties of the CSC-IPCA estimator. The numbers of treated units and post-treatment periods are fixed at $N_{treat}=5$ and $T_{post}=5$. We vary the number of control units $N_{ctrl}$, pre-treatment periods $T_{pre}$, and the proportion of observed covariates $\alpha$ with the total number of covariates $L=9$ to investigate the finite sample properties. As shown in Table 1, the bias, RMSE, and STD are estimated based on 1000 simulations [^9] The root mean squared error (RMSE) is defined as $RMSE=\sqrt{\frac{1}{T_{pre}}\sum_{t\in T_{pre}}\left(ATT_{t}-\widehat{ATT}_{t}\right)^{2}}$. The standard deviation (STD) is defined as $STD=\frac{1}{T_{pre}}\sum_{t\in T_{pre}}\left(\widehat{ATT}_{t}-\frac{1}{T_{pre}}\sum_{t\in T_{pre}}\widehat{ATT}_{t}\right)^{2}$. The results indicate that the bias of the CSC-IPCA estimator decreases as the number of control units and pre-treatment periods increases. As suggested by the theoretical results, the convergence rate of the CSC-IPCA estimator is the smaller one between $\mathcal{O}_{p}\left(\sqrt{N_{ctrl}}\right)$ and $\mathcal{O}_{p}\left(\sqrt{N_{treat}T_{pre}}\right)$. In this simulation example, since the smaller one is always $\mathcal{O}_{p}\left(\sqrt{N_{ctrl}}\right)$, we observe that the bias decreases significantly when the number of control units increases from 10 to 40. The number of observed covariates also plays a crucial role in bias reduction. The bias decreases the most when the proportion of observed covariates increases from $1/3$ to 1 (all covariates are observed). We observe a similar pattern in RMSE and STD. It is worth noting that if we observe all the covariates (i.e., $\alpha=1$), the bias, RMSE, and STD of the CSC-IPCA estimator all reduce to the lowest levels, even with a small number of control units and pre-treatment periods. We present the finite sample properties of the CSC-IFE and SCM estimators in Appendix C.2. When observing all the covariates, the CSC-IFE estimator is comparable to the CSC-IPCA estimator; however, when the number of observed covariates decreases, the CSC-IFE estimator becomes more biased than the CSC-IPCA estimator. The SCM estimator is always biased due to the settings.

### v2

@@ 3.2 Assumptions for identification
To ensure the identification of the treatment effect, we introduce a series of assumptions about the DGPs, in addition to the assumption regarding the functional form. These assumptions are crucial for achieving consistent estimation of the ATT, yet they are easy to satisfy in practice.
@@
    Covariate orthogonality: $\mathbb{E}\left[X^{\prime}_{it}\epsilon_{it}\right]=\textbf{0}_{L\times 1}$,
@@
    Moment condition of factors and covariates. The following moments exist: $\mathbb{E}\|F_{t}F^{\prime}_{t}\|^{2}$, $\mathbb{E}\|X^{\prime}_{it}\epsilon_{it}\|^{2}$, $\mathbb{E}\|X^{\prime}_{it}X_{it}\|^{2}$, $\mathbb{E}\left[\|X^{\prime}_{it}X_{it}\|^{2}\|F_{t}F^{\prime}_{t}\|^{2}\right]$,
@@
    Almost surely, $X_{it}$ is bounded, and define $\Omega_{t}^{xx}:=\mathbb{E}\left[X_{it}^{\prime}X_{it}\right]$, then almost surely, $\Omega_{t}^{xx}>\epsilon$ for some $\epsilon>0$.
  4. (4)
The assumptions outlined above serve as regularity conditions necessary for ensuring the consistency of the estimator. The first condition reflects the exclusion restriction commonly associated with instrumental variable regression, indicating that the instruments (in this case, covariates) are orthogonal to the error terms. This orthogonality ensures that the covariates do not correlate with any unobserved error terms influencing the outcome. Similarly, the functional form described in Equation 4 aligns with the framework utilized in two-stage instrumental regression. The second condition imposes a limitation on the second moment of a series of random variables, ensuring that their variances remain finite and do not escalate indefinitely. The third condition mandates that the parameter space for the mapping matrix $\Gamma$ to be compact, thereby circumventing issues related to rank deficiency. This compactness is vital for the estimability of $\Gamma$ and the existence of its inverse, mirroring standard factor analysis assumptions.
To derive the asymptotic properties of the CSC-IPCA estimator, we need to introduce additional assumptions. Assumption 3 encompasses panel-wise and cross-sectional central limit theorems for various variables, which are fulfilled by diverse mixing processes. These conditions are pivotal for determining the asymptotic distribution of common factors $F_{t}$ and mapping matrix $\Gamma$ estimations. For an in-depth discussion, we refer to Kelly et al. (2020). The requirement of bounded dependence stipulates that both time series and cross-sectional dependencies of $X_{it}\epsilon_{it}$ are bounded, a crucial step for establishing asymptotic normality. Meanwhile, the assumption of cross-sectional homoskedasticity simplifies the expressions for the asymptotic variances of the estimated mapping matrix $\Gamma$. It should be noted that relaxing this assumption would not alter the convergence rate.
@@
Assumptions for asymptotic normality:
@@
[…]
@@
[…]
  4. (4)
    Bounded dependence: $\frac{1}{NT}\sum_{i,j,t,s}\|\tau_{ij,ts}\|<\infty$, where $\tau_{ij,ts}:=\mathbb{E}\left[X^{\prime}_{it}\epsilon_{it}\epsilon^{\prime}_{js}X_{js}\right]$
  5. (5)
@@ 4 Estimation
The CSC-IPCA estimator of the treatment effect for a treated unit $i$ at time $t$ is defined as the difference between the observed outcome and its imputed counterfactual: $\hat{\delta}_{it}=Y_{it}^{1}-\hat{Y}_{it}^{0}$. To combine the functional form in Equation 4, we get the structural component of the DGPs for outcome $Y_{it}=(X_{it}\Gamma)F^{\prime}_{t}$. The CSC-IPCA method is estimated by minimizing the sum of squared residuals of the following objective function:
@@
Unlike the IFE method (Bai (2009); Xu (2017)), our approach requires estimating only two parameters, $\Gamma$ and $F_{t}$, simplifying the process. Different from principal component analysis (Jolliffe (2002); Stock and Watson (2002)), our method involves using covariates to instrument the factor loadings. This necessitates the estimation of $\Gamma$ rather than $\Lambda_{i}$, so we can not directly use eigenvalue decomposition. While the objective function in Equation 7 formulates the problem as minimizing a quadratic function with a single unknown variable (e.g., $\Gamma$) while holding the other variable (e.g., $F_{t}$) constant. This structure enables the application of the alternating least squares (ALS) method for a numerical solution. Generally, the imputation for the missing counterfactual $\hat{Y}_{it}^{0}$ is executed in four steps:
@@
Step 2: The subsequent step involves estimating the mapping matrix $\hat{\Gamma}_{treat}$ for treated unit $i$ at time $t$, employing the previously estimated time-varying factors $\hat{F}_{t}$, using treated group data exclusively for the pre-treatment period.
@@
$F_{norm,t}$ $=R^{-1}\hat{F}_{t},$
$s.t.\Gamma_{norm}^{\prime}\Gamma_{norm}$ $=\mathcal{I}_{K},\quad F_{norm,t}F_{norm,t}^{\prime}/T=\text{Diagonal}.$
Similar to most factor analysis methods, the estimated $\Gamma$ and $F_{t}$ are not deterministic. There exists an infinite number of “rotated” parameters $\Gamma R$ and $R^{-1}F_{t}$ that yield the same objective function value (i.e., $X_{it}\Gamma F_{t}=X_{it}\Gamma RR^{-1}F_{t}$). To make the estimation identifiable, it is necessary to impose some constraints on the estimated $\Gamma$ and $F_{t}$. Following Connor and Korajczyk (1993); Stock and Watson (2002); Bai and Ng (2002), the aforementioned restrictions reduce the model’s complexity and make it easier to understand and interpret the relationships between the factors $F_{t}$ and factor loadings $\Lambda_{it}$. Unlike the estimation methods in Bai (2009); Xu (2017), where normalization constraints are set before the estimation, the structural component $X_{it}\Gamma F_{t}$ allows us to normalize the estimated $\Gamma$ and $F_{t}$ after the estimation. This enables us to easily find the rotation matrix $R$ that satisfies the above constraints.
Step 4: The final step involves imputing the counterfactual outcome $\hat{Y}_{it}^{0}$ for treated unit $i$ at time $t$ by substituting the estimated mapping matrix $\hat{\Gamma}_{norm}$ and the common factors $\hat{F}_{norm,t}$ into the following equation:
$\hat{Y}_{it}(0)=(X_{it}\hat{\Gamma}_{norm})\hat{F}^{\prime}_{norm,t},\quad\forall i\in\mathcal{T}\ \&\ T_{pre}<t\leq T.$ (11)
The main difference between CSC-IPCA and the instrumented principal component analysis (IPCA) as proposed by Kelly et al. (2020) lies in the purpose of prediction. In the IPCA method, the authors predict the next period stock returns using all covariates from the preceding period, under the assumption that the mapping matrix $\Gamma$ remains constant across all observations. In contrast, CSC-IPCA introduces a pivotal distinction: it operates under the assumption that treated and control groups are characterized by unique mapping matrices, $\Gamma_{\text{treat}}$ and $\Gamma_{\text{ctrl}}$. This assumption is vital for the unbiased estimation of the ATT, setting the CSC-IPCA method apart by directly addressing heterogeneity in treated and control groups (also result in who receives the treatment) through the specification of group-specific mapping matrices. The detailed estimation procedures are presented in the Appendix A.2.
@@ 4.2 Inference
In the context of causal inference, the pioneering application of factor models is attributed to Hsiao et al. (2012). Nonetheless, a formal framework for inference using this method was not established until the works of Chan et al. (2016) and Li (2018). These methods of inference rely on a large number of control units and pre-treatment periods to develop asymptotic properties. In recent years, conformal inference developed by Chernozhukov et al. (2021) has gained popularity in the causal inference literature, as seen in Ben-Michael et al. (2021), Roth et al. (2023), and Imbens (2024). Conformal inference is a nonparametric method that provides exact and robust inference without requiring the specification of the model. Our causal inference framework, designed for predicting missing counterfactuals, allows us to construct inference procedures based on conformal prediction, as introduced by Shafer and Vovk (2008), to ensure robustness against misspecification. The causal effect is identified as the difference between the observed outcomes and these estimated counterfactuals, expressed mathematically as:
$\theta_{it}=Y^{1}_{it}-\hat{Y}^{0}_{it},\quad\forall i\in N_{treat}\ \&\ T_{pre}<t\leq T,$
where $\theta_{it}$ denotes the treatment effect for unit $i$ at time $t$, and $\hat{Y}^{0}_{it}$ represents the imputed counterfactual outcome.
@@ 5 Monte Carlo Simulation
We initiate our analysis with a DGP that incorporates $L=10$ covariates and $K=3$ common factors, along with unit and time-fixed effects:
@@
where $X_{it}=[x_{it}^{1},\ldots,x_{it}^{L}]$ denotes a vector of $L\times 1$ covariates, which follows a VAR(1) process. $X_{it}=\mu_{i}+A_{i}X_{i,t-1}+\nu_{it}$, where $A_{i}$ is a $L\times L$ variance-covariance matrix [^4] In our methodology, the variance-covariance matrix is not constrained to be diagonal, thus allowing covariates within each unit to be correlated, reflecting the typical scenario in most economic time series data. To emphasize the independence among different units, we generate $N$ unique variance-covariance matrices, each corresponding to a unit, ensuring cross-sectional independence and preserving time-series correlation. Moreover, we impose a condition on these matrices by requiring the eigenvalues of $A_{i}$ to have characteristic roots that reside inside the unit circle, thereby assuring the stationarity of the VAR(1) process., The drift term $\mu_{i}$ equals 0 for control units and 2 for treated units [^5] This configuration underscores that the treatment assignment is not random; rather, it depends on the covariates $X_{it}$., and $\nu_{it}$ is a $L\times 1$ vector of i.i.d. standard normal errors. While $F_{t}=[f_{t}^{1},\ldots,f_{t}^{K}]$ denotes the vector of common factors, adhering to a similar VAR(1) process, the variable $\epsilon_{it}$ represents the idiosyncratic error term. Unit and time fixed effects, $\alpha_{i}$ and $\xi_{t}$ respectively, are uniformly drawn from the interval $(0,1)$. The coefficient vector $\beta=[\beta^{1},\ldots,\beta^{L}]$ associated with the covariates is drawn uniformly from $(0,1)$, and $\Gamma$, the $L\times K$ mapping matrix, is drawn uniformly from $(-0.1,0.1)$. The treatment indicator $D_{it}$ is binary, defined as $D_{it}=1$ for treated units during post-treatment periods, and $D_{it}=0$ otherwise. The heterogeneous treatment effect is modeled as $\delta_{it}=\bar{\delta}_{it}+e_{it}$, where $e_{it}$ is i.i.d as standard normal, and $\bar{\delta_{t}}=[0,\cdots,0,1,2,\ldots,T_{post}]$ represents a time-varying treatment effect [^6] Here we simplify the treatment effect to be constant across units, however the heterogeneous treatment effect across units can also be easily employed.. Only the outcome $Y_{it}$, the covariates $X_{it}$, and the treatment indicator $D_{it}$ are observed, while all other variables remain unobserved.
@@
In this graphic, the upper panel plots simulated data following the above DGPs. The light blue lines represent treated units and the light gray lines represent controls. Key parameters are $N_{treat}=5,N_{ctrl}=45,T_{0}=20,T_{1}=10,L=10$. The lower panel plots a simple event study.
Figure 1 represents the simulated data following our data generating process. Observations from the upper panel indicate that the parallel trend assumption is not met. To verify this, we plot a simple event study, which clearly revealing a failure in the parallel trend assumption. Furthermore, outcomes for treated units are marginally higher than for control units. In such cases, the synthetic control method will be biased, as it avoids extrapolation and typically fits poorly for treated units.
@@ 5.1 A simulated example
Following this DGPs, Figure 2 illustrates both the raw data and the imputed counterfactual outcomes as estimated by the CSC-IPCA method. In the upper panel, control units are represented in gray and treated units in light blue, with the average outcome for treated units highlighted in orange. The imputed synthetic average for treated outcomes is also shown, delineated by an orange dashed line. The CSC-IPCA method is capable of capturing the trajectory of the average outcome for treated units before treatment, and we observe the divergence after the treatment. The lower panel of Figure 2 shows the estimated ATT (dashed line) alongside the true ATT (solid line) and the 95% confidence interval based on conformal inference. The CSC-IPCA method is able to capture the true ATT, as evidenced by the close alignment between the dashed and solid lines. The confidence interval constructed through conformal inference is also accurate, as it encompasses the true ATT.
@@ 5.3 Finite sample properties
We present the Monte Carlo simulation results in Table 1 to investigate the finite sample properties of the CSC-IPCA estimator. The numbers of treated units and post-treatment periods are fixed at $N_{treat}=5$ and $T_{post}=5$. We vary the number of control units $N_{ctrl}$, pre-treatment periods $T_{pre}$, and the proportion of observed covariates $\alpha$ with the total number of covariates $L=9$ to investigate the finite sample properties. As shown in Table 1, the bias, RMSE, and STD are estimated based on 1000 simulations [^7] The root mean squared error (RMSE) is defined as $RMSE=\sqrt{\frac{1}{T_{pre}}\sum_{t\in T_{pre}}\left(ATT_{t}-\widehat{ATT}_{t}\right)^{2}}$. The standard deviation (STD) is defined as $STD=\frac{1}{T_{pre}}\sum_{t\in T_{pre}}\left(\widehat{ATT}_{t}-\frac{1}{T_{pre}}\sum_{t\in T_{pre}}\widehat{ATT}_{t}\right)^{2}$. The results indicate that the bias of the CSC-IPCA estimator decreases as the lower bound, which is the smaller one between the number of control units and pre-treatment periods, increases. As suggested by the theoretical results, the convergence rate of the CSC-IPCA estimator is the smaller one between $\mathcal{O}_{p}\left(\sqrt{N_{ctrl}}\right)$ and $\mathcal{O}_{p}\left(\sqrt{N_{treat}T_{pre}}\right)$.
In this simulation example, since the smaller one is always $\mathcal{O}_{p}\left(\sqrt{N_{ctrl}}\right)$, we observe that the bias decreases significantly when the number of control units increases from 10 to 40. The number of observed covariates also plays a crucial role in bias reduction. The bias decreases the most when the proportion of observed covariates increases from $1/3$ to 1 (all covariates are observed). We observe a similar pattern in RMSE and STD. It is worth noting that if we observe all the covariates (i.e., $\alpha=1$), the bias, RMSE, and STD of the CSC-IPCA estimator all reduce to the lowest levels, even with a small number of control units and pre-treatment periods. We present the finite sample properties of the CSC-IFE and SCM estimators in Appendix C.2. When observing all the covariates, the CSC-IFE estimator is comparable to the CSC-IPCA estimator; however, when the number of observed covariates decreases, the CSC-IFE estimator becomes more biased than the CSC-IPCA estimator. The SCM estimator is always biased due to the settings.